It costs Homers Manufacturing $0.45 to produce baseballs and Homer sells them for $5 a piece. Homer pays a sales commissionLucas Industries uses departmental overhead rates to allocate its manufacturing overhead to jobs. The company has two departmIt costs Homers Manufacturing $0.65 to produce baseballs and Homer sells them for $7.00 a piece. Homer pays a sales commissi
Answer #1

1. The right option is "B. $294,550".

Contribution margin = Sales - Variable expenses

Where, sales = Selling price per unit * 68,500 baseballs = $5 * 68,500 = $342,500

Variable expenses = ( Manufacturing costs per unit + Sales commission per unit ) * 68,500 baseballs = ( $0.45 per unit + 5% of $5 ) * 68,500 = $47,950

Contribution margin = Sales - Variable expenses = $342,500 - $47,950 = $294,550.

2. The right option is "B. $510".

Manufacturing overhead allocated to JOB 603 using departmental overhead rates

Total manufacturing overhead = Manufacturing overhead in Assembly department + Manufacturing overhead in Sanding department

where, Manufacturing overhead in Assembly department = Overhead rate per machine hours * Machine hours used = $40 * 11 = $440

Manufacturing overhead in Sanding department = Overhead rate per direct labor hours * Direct labor hours used = $10 * 7 = $70.

Total manufacturing overhead = Manufacturing overhead in Assembly department + Manufacturing overhead in Sanding department = $440 + $70 = $510.

3. The right option is " C. $428,625".

Gross profit = Sales - Cost of goods sold

where, sales = Selling price per unit * 67,500 units = $7 * 67,500 units = $472,500.

Cost of goods sold = manufacturing cost per unit * 67,500 units = $0.65 per unit * 67,500 units = $43,875

Gross profit = Sales - Cost of goods sold = $472,500 - $43,875 = $428,625.